40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 23 01:27:11 2013, Manasa, Summary, LSC, Xarm beat note search continues 
    Reply  Thu May 23 13:14:34 2013, Koji, Summary, LSC, Xarm beat note search continues 
Message ID: 8624     Entry time: Thu May 23 01:27:11 2013     Reply to this: 8629
Author: Manasa 
Type: Summary 
Category: LSC 
Subject: Xarm beat note search continues 

Towards finding the x-arm beat note:

The green would not lock to a maximum GTRX this morning. In the course of aligning the green stably to the X arm, somewhere down the line, the input pointing got messed up (reasons unknown). To set this right, Koji tried to lock the Yarm with POY DC but it wouldn't work. The transmon for Y had to be set up temporarily and the Y arm was locked with TRY. This restored the input pointing and the arms locked with transmission TRX/TRY > 0.9 counts. The transmon path along the Y arm was then re-configured as mentioned in Annalisa's elog.

I still had trouble getting the X-green locked in TEM00 (similar situation mentioned by Jenne in elog). The arm cavity mirrors were tweaked to get the green to resonate in TEM00 but it wouldn't stay locked when the temperature of the x-end NPRO was changed. Koji helped recover missing links to filters for the ALS_X_SLOW servo from the archives. Enabling the filters helped keep the green locking stable for laser temperature changes (which corresponds to 'offset' change in ALS_X_SLOW servo screen).

PSL green alignment was checked once again and the X-end laser temperature was scanned trying to find the beatnote. RFMON from the beatbox was connected to the spectrum analyzer. I have scanned through the whole range of offset but have not been able to find the beat note yet.


The search will continue tomorrow

ELOG V3.1.3-